OpenStudy (anonymous):

Enter the prime factorization of 154

ganeshie8 (ganeshie8):

154 = 2*77

ganeshie8 (ganeshie8):

154 = 2*77 = 2*11*7

ganeshie8 (ganeshie8):

are 2, 11, 7 are prime ?